# Bias–variance Tradeoff in Tensor Estimation

### v1

@@ Abstract
We study denoising of a third-order tensor when the ground-truth tensor is not necessarily Tucker low-rank. Specifically, we observe
@@
where $X^{\ast}$ is the ground-truth tensor, and $Z$ is the noise tensor. We propose a simple variant of the higher-order tensor SVD estimator $\widetilde{X}$. We show that uniformly over all user-specified Tucker ranks $(r_{1},r_{2},r_{3})$,
$\|\widetilde{X}-X^{*}\|_{\mathrm{F}}^{2}=O\Big(\kappa^{2}\Big\{r_{1}r_{2}r_{3}+\sum_{k=1}^{3}p_{k}r_{k}\Big\}\;+\;\xi_{(r_{1},r_{2},r_{3})}^{2}\Big)\quad\text{ with high probability.}$
Here, the bias term $\xi_{(r_{1},r_{2},r_{3})}$ corresponds to the best achievable approximation error of $X^{\ast}$ over the class of tensors with Tucker ranks $(r_{1},r_{2},r_{3})$; $\kappa^{2}$ quantifies the noise level; and the variance term $\kappa^{2}\{r_{1}r_{2}r_{3}+\sum_{k=1}^{3}p_{k}r_{k}\}$ scales with the effective number of free parameters in the estimator $\widetilde{X}$. Our analysis achieves a clean rank-adaptive bias–variance tradeoff: as we increase the ranks of estimator $\widetilde{X}$, the bias $\xi(r_{1},r_{2},r_{3})$ decreases and the variance increases. As a byproduct we also obtain a convenient bias-variance decomposition for the vanilla low-rank SVD matrix estimators.
@@ 1 Introduction
Low-rank tensor models are a cornerstone of modern machine learning. They capture essential structure in high-dimensional signals while offering substantial gains in computation and storage. Applications include recommender systems (Koren et al.,, 2009), topic modeling (Blei et al.,, 2003), community detection (Anandkumar et al.,, 2014; Abbe,, 2018), and, more recently, latent variable learning (Diakonikolas and Kane,, 2024) and generative modeling (Hur et al.,, 2023; Peng et al.,, 2023). The effectiveness of these methods demonstrate that many real-world signals embedded in high-dimensional ambient spaces can be well-approximated by low-rank structures.
@@
The study of exactly low-rank estimation is well-developed, particularly in the matrix setting. A large body of work has introduced efficient algorithms for tasks such as matrix completion and denoising, including nuclear norm relaxation (Koltchinskii et al.,, 2011), convex optimization (Candes and Recht,, 2012), nonconvex methods (Chi et al.,, 2019), and SVD guarantees with unbalanced matrices (Cai and Zhang,, 2018).
@@
Parallel efforts in low-rank tensor estimation are well-developed, including the study of higher-order tensor singular value decompositions in the non-random setting (De Lathauwer et al., 2000a, ; De Lathauwer et al., 2000b, ) and Riemannian optimization methods for low-rank Tucker tensor completion (Kressner et al.,, 2014). Nevertheless, the tensor setting poses additional challenges: many fundamental problems are computationally intractable, multiple inequivalent notions of rank (e.g. CP and Tucker) exist, and best low-rank approximations need not be unique (e.g. Hillar and Lim,, 2013). Most theoretical guarantees in tensor estimation (e.g. Zhang and Xia,, 2018; Han et al.,, 2022), like their matrix counterparts, rely on the assumption that the underlying signal is exactly low-rank.
In practice, however, signals are not exactly low-rank: their singular spectra typically decay gradually rather than vanishing at a finite rank. This creates an interesting gap between theory and practice, since existing analyses provide guarantees under idealized exact low-rankness and leave open the question of how to analyze estimation error in more realistic settings.
We address this gap by designing and analyzing low-rank tensor estimation when the signal is not low-rank. In such settings, estimation error is governed by a bias–variance tradeoff: truncating to a smaller rank incurs approximation bias, while larger ranks inflate variance by amplifying noise. In particular, we propose a simple variant of the higher-order singular value decomposition (HOSVD) algorithm and provide the analysis of tensor estimation without assuming low-rankness. Our main result (Theorem 1) establishes an explicit bias–variance decomposition of the estimation error:
  • The bias term quantifies the approximation error incurred by truncating a spectrum.
  • The variance term matches the optimal rate for exactly low-rank models.
The success of SVD and HOSVD has been well established by their decades of widespread use. The objective of our manuscript is to provide mathematical justification for the information-efficient bias-variance tradeoff achieved by these two classical algorithms, and to establish uniform performance guarantees that hold for any user-specified ranks.
To achieve this clean bias-variance tradeoff, we utilize classical linear algebra results such as Mirsky’s and Ky Fan’s theorems to develop non-trivial extensions of classical perturbation tools to general tensors. As a convenient side result, we also recover a clean bias–variance characterization for matrices, thereby unifying the matrix and tensor perspectives. Finally, we validate our theory with simulations that confirm the existence of tradeoffs across different regimes of spectral decay and noise.
@@ 1.1 Notation
2 Bias–variance tradeoff in tensor estimation
@@
where $X^{*}$ is an unknown signal and $Z$ is a random perturbation. Even when $X^{*}$ has a full Tucker rank $(p_{1},p_{2},p_{3})$, it is common to approximate $X^{*}$ by a Tucker rank-$(r_{1},r_{2},r_{3})$ estimator. This is because (i) low-rank structure yields dramatic gains in computation and memory, see Remark 2; and (ii) the approximation bias can decay quickly when approximating $X^{*}$ by a Tucker low-rank tensor, so balancing bias and variance may lead to a smaller estimation error. Given a target Tucker rank $(r_{1},r_{2},r_{3})$, we consider the following simple variant of the HOSVD algorithm (De Lathauwer et al., 2000b, ).
@@
$\mathcal{T}_{(r_{1},r_{2},r_{3})}=\big\{A\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}:\ \operatorname*{rank}(\mathcal{M}_{k}(A))\leq r_{k},\ k=1,2,3\big\}.$
@@
Tightness. Our upper bound in (3) attains the minimax optimal rate (up to constants) over the class of tensors with Tucker rank at most $(r_{1},r_{2},r_{3})$. Indeed, by definition, $\xi_{(r_{1},r_{2},r_{3})}$ is the smallest possible approximation error of $X^{*}$ over $\mathcal{T}_{(r_{1},r_{2},r_{3})}$. In addition, Zhang and Xia, (2018) prove that, when $X^{*}$ is exactly low-rank with Tucker rank $(r_{1},r_{2},r_{3})$, the term $\kappa^{2}\Bigl(\sum_{k=1}^{3}p_{k}r_{k}+r_{1}r_{2}r_{3}\Bigr)$ is minimax optimal up to a universal constant. For arbitrary (e.g. full-rank) tensors, our bound should be interpreted as an upper bound; a matching minimax lower bound for that broader class is not known.
Proof techniques. For the bias, we leverage classical linear algebra tools, such as Mirsky’s and Ky Fan’s theorems, to characterize the optimal approximation error appearing in (3). To handle the variance of the estimator $\widetilde{X}$, we adopt existing techniques from Zhang and Xia, (2018).
@@
In Theorem 1, we assume that the entries of the noise tensor $Z$ is i.i.d. sub-Gaussian. This is a commonly seen condition in the tensor literature (e.g. Zhang and Xia,, 2018; Han et al.,, 2022).
The signal-to-noise ratio (SNR) condition (2) in Theorem 1 is also commonly seen in literature. In fact, if the ground truth $X^{*}$ is exactly Tucker low-rank with Tucker rank $(r_{1},r_{2},r_{3})$, then $\sigma_{r_{k}+1}(\mathcal{M}_{k}(X^{*}))=0$ and (2) reduces to the SNR condition used in Zhang and Xia, (2018), when $r_{\max}$ is a bounded constant.
Remark 2.
@@
$Y_{(r_{1},r_{2},r_{3})}=S\times_{1}U_{1}\times_{2}U_{2}\times_{3}U_{3},$
@@
$s=t_{2}^{\top}S^{\prime}\in\mathbb{R}^{r_{3}}\quad(\text{cost }\approx 2r_{2}r_{3}),\quad w=s^{\top}t_{3}\in\mathbb{R}\quad(\text{cost }\approx r_{3}).$

### v3

# Optimal Bias-variance Tradeoff in Matrix and Tensor Estimation

@@ Abstract
We study matrix and tensor denoising when the underlying signal is not necessarily low-rank. In the tensor setting, we observe
@@
where $X^{\ast}$ is an unknown signal tensor and $Z$ is a noise tensor. We propose a one-step variant of the higher-order SVD (HOSVD) estimator, denoted $\widetilde{X}$, and show that, uniformly over any user-specified Tucker ranks $(r_{1},r_{2},r_{3})$, with high probability,
$\|\widetilde{X}-X^{\ast}\|_{\mathrm{F}}^{2}=O\Big(\kappa^{2}\Big\{r_{1}r_{2}r_{3}+\sum_{k=1}^{3}p_{k}r_{k}\Big\}+\xi_{(r_{1},r_{2},r_{3})}^{2}\Big).$
Here, $\xi_{(r_{1},r_{2},r_{3})}$ is the best achievable Tucker rank-$(r_{1},r_{2},r_{3})$ approximation error of $X^{\ast}$ (bias), $\kappa^{2}$ quantifies the noise level, and $\kappa^{2}\{r_{1}r_{2}r_{3}+\sum_{k=1}^{3}p_{k}r_{k}\}$ is the variance term scaling with the effective degrees of freedom of $\widetilde{X}$. This yields a rank-adaptive bias–variance tradeoff: increasing $(r_{1},r_{2},r_{3})$ decreases the bias $\xi_{(r_{1},r_{2},r_{3})}$ while increasing variance. In the matrix setting, we show that truncated SVD achieves an analogous bias–variance tradeoff for arbitrary signal matrices. Notably, our matrix result requires no assumptions on the signal matrix, such as finite rank or spectral gaps. Finally, we complement our upper bounds with matching information-theoretic lower bounds, showing that the resulting bias–variance tradeoff is minimax optimal up to universal constants in both the matrix and tensor settings.
@@ 1 Introduction
Low-rank matrix and tensor estimation is a cornerstone of modern machine learning. Such models capture essential structure in high-dimensional signals while offering substantial gains in computation and storage. Applications include recommender systems (Koren et al., 2009), topic modeling (Blei et al., 2003), community detection (Abbe, 2018; Diakonikolas and Kane, 2024), and, more recently, latent variable learning (Anandkumar et al., 2014; Sherman and Kolda, 2020; Zhang and Kileel, 2023) and generative modeling (Hur et al., 2023; Peng et al., 2023). The success of these methods suggests that many real-world signals in high-dimensional ambient spaces can be well approximated by low-rank structure.
@@
The theory of low-rank estimation is well developed in the matrix setting. Many efficient methods have been proposed for matrix completion and denoising, including nuclear-norm relaxation (Koltchinskii et al., 2011), convex optimization (Candes and Recht, 2012), nonconvex approaches (Chi et al., 2019), and refined SVD guarantees for highly rectangular matrices (Cai and Zhang, 2018). Nevertheless, most existing guarantees are tailored to exactly low-rank signals. When the signal is not exactly low-rank, accurate estimation is typically established under additional structural conditions, such as spectral gap assumptions.
@@
Substantial efforts have advanced low-rank tensor estimation, including higher-order tensor singular value decompositions in deterministic settings (De Lathauwer et al., 2000a, b), symmetric tensor decomposition in deterministic settings (Jin et al., 2024), and Riemannian optimization methods for low-rank Tucker tensor completion (Kressner et al., 2014). At the same time, the tensor setting poses additional challenges: many fundamental problems are computationally intractable, multiple inequivalent notions of rank (e.g., CP and Tucker) coexist, and best low-rank approximations need not be unique (e.g. Hillar and Lim, 2013). Accordingly, most theoretical guarantees in tensor estimation (e.g. Anandkumar et al., 2017; Zhang and Xia, 2018; Han et al., 2022), like their matrix counterparts, assume that the underlying signal is exactly low-rank. In practice, however, signals are rarely exactly low-rank: their singular values typically decay gradually rather than dropping to zero at a finite rank. This creates a gap between theory and practice, since many existing guarantees are tailored to idealized exactly low-rank models and do not directly characterize performance in realistic regimes.
We close this gap with a rank-adaptive analysis of classical spectral estimators in the approximately low-rank regime. Our contributions are as follows:
  • Rank-adaptive bias–variance theory. We establish an explicit bias–variance decomposition for the estimation error that holds uniformly over user-specified target ranks.
  • Matrix setting. We show that truncated SVD achieves the optimal bias–variance tradeoff without any structural assumptions on the signal matrix.
  • Tensor setting. We analyze a simple one-step variant of HOSVD and obtain analogous guarantees without assuming that the ground truth is exactly Tucker low-rank.
  • Minimax optimality. We complement our upper bounds with matching information-theoretic lower bounds, establishing minimax optimality up to universal constants.
  • Methodology. Our analysis combines classical linear algebra tools, including Mirsky’s and Ky Fan’s theorems, with modern concentration arguments. As a byproduct, we develop new matrix and tensor perturbation bounds that may be of independent interest.
  • Numerical validation. We validate the predicted tradeoffs through simulations and a real 3D brain MRI example.
@@ 1.1 Notation
2 Optimal bias–variance tradeoff in matrix estimation
We consider the model
$Y=X^{*}+Z\in\mathbb{R}^{m\times n},$ (1)
where $X^{*}$ is an unknown population matrix with arbitrary rank, and $Z$ is a noise matrix. In particular, we do not assume that $X^{*}$ is low-rank, nor do we assume any spectral gap conditions on $X^{*}$. Even when $\mathrm{rank}(X^{*})$ is not small, it is common to approximate $X^{*}$ by a rank-$r$ estimator, because (i) low-rank decomposition can dramatically increase the computational and memory efficiency, see Remark 1 for more details; and (ii) the approximation bias can be small, so balancing the bias and variance may lead to smaller estimation error. Let
By the Eckart–Young–Mirsky theorem (Eckart and Young, 1936), $X_{(r)}^{*}$ is the best rank-$r$ approximation to $X^{*}$ in Frobenius norm in the sense that
Suppose we observe $Y$ instead of $X^{*}$, and write the SVD of $Y$ as
Theorem 1.
Under (1), let $Y_{(r)}$ be defined as in (2). Then
Assumption-free with respect to $X^{*}$. Theorem 1 requires no structural conditions on $X^{*}$. In particular, $\operatorname*{rank}(X^{*})$ may be arbitrary, and no spectral gap assumptions are needed.
In what follows, we illustrate the applicability and optimality of Theorem 1 in two popular matrix estimation settings.
Corollary 1.
$\|v^{\top}Zw\|_{\psi_{2}}\leq\kappa.$
[…]
Next, we show that the error bound in Corollary 1 is minimax optimal.
Theorem 2.
Let $r,m,n$ be arbitrary positive integers such that $r\leq\min\{m,n\}$. For $\xi>0$, let $\mathcal{M}_{r}^{\xi}$ be such that
[…]
Consider the model $Y=X^{\ast}+Z\in\mathbb{R}^{m\times n}$ where the entries of $Z$ are i.i.d. $\mathrm{subGaussian}(0,\kappa^{2})$. Suppose in addition that $\xi^{2}\leq(m-r)n\kappa^{2}$. Then there exist a universal constant $c>0$ such that
$\inf_{\widehat{X}}\sup_{X^{\ast}\in\mathcal{M}_{r}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{\ast}\bigr\|_{\mathrm{F}}^{2}\$$\geq\ c\Bigl\{\kappa^{2}r(m+n)+\xi^{2}\Bigr\}.$
Corollary 1 and Theorem 2 together show that (3) is minimax optimal when $Z$ is sub-Gaussian. Next, we illustrate the applicability of Theorem 1 in covariance matrix estimation.
Corollary 2.
$Y=\frac{1}{N}\sum_{k=1}^{N}Z_{k}Z_{k}^{\top},$
$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq C\Bigl(\kappa\sqrt{r}\Bigl[\sqrt{\frac{n}{N}}+\frac{n}{N}\Bigr]+\xi_{(r)}\Bigr),$
Remark 1 (Computational and storage savings from matrix low-rank factorization).
$t=V_{(r)}^{\top}v\in\mathbb{R}^{r}\ (\text{cost}\approx 2nr),\quad s=\Sigma_{(r)}t\in\mathbb{R}^{r}\ (\text{cost}\approx r),\quad w=U_{(r)}s\in\mathbb{R}^{m}\ (\text{cost}\approx 2mr).$
The SVD factorization also reduces storage costs. Storing the full matrix $Y$ requires $O(mn)$ memory, whereas storing the factors $U_{(r)}\in\mathbb{R}^{m\times r}$, $\Sigma_{(r)}\in\mathbb{R}^{r\times r}$, and $V_{(r)}\in\mathbb{R}^{n\times r}$ requires only $O(mr+nr)$ scalars.
3 Optimal bias–variance tradeoff in tensor estimation
@@
where $X^{*}$ is an unknown signal tensor and $Z$ is a noise tensor. Importantly, we note that the results in this section do not assume that $X^{*}$ is exactly Tucker low-rank. Given a user-specified Tucker rank $(r_{1},r_{2},r_{3})$, we consider the following simple variant of the HOSVD algorithm (De Lathauwer et al., 2000b).
@@
$\mathcal{T}_{(r_{1},r_{2},r_{3})}=\big\{A\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}:\ \operatorname*{rank}(\mathcal{M}_{k}(A))\leq r_{k},\ k=1,2,3\big\}.$ (5)
@@
Proof techniques. For the bias, we leverage classical linear algebra tools, such as Mirsky’s and Ky Fan’s theorems, to characterize the optimal approximation error appearing in (7). To handle the variance of the estimator $\widetilde{X}$, we adopt existing techniques from Zhang and Xia (2018).
@@
In Theorem 3, we assume that the entries of the noise tensor $Z$ are i.i.d. sub-Gaussian. This is a commonly seen condition in the tensor literature (e.g. Zhang and Xia, 2018; Han et al., 2022).
The signal-to-noise ratio (SNR) condition (6) in Theorem 3 is also commonly seen in literature. In fact, suppose the ground truth $X^{*}$ is exactly Tucker low-rank with Tucker rank $(r_{1},r_{2},r_{3})$. Then $\sigma_{r_{k}+1}(\mathcal{M}_{k}(X^{*}))=0$, and (6) reduces to the SNR condition used in Zhang and Xia (2018) when $r_{\max}$ is a bounded constant.
We present a relative Frobenius error bound as a corollary of Theorem 3.
Corollary 3 (Relative estimation error).
Suppose the conditions of Theorem 3 hold and $r_{k}\leq\sqrt{p_{\min}}$ for $k=1,2,3$. Then, with probability at least $1-C_{1}p_{1}p_{2}p_{3}\exp(-C_{2}p_{\min})$,
$\frac{\|\widetilde{X}-X^{*}\|_{\mathrm{F}}}{\|X^{*}\|_{\mathrm{F}}}\;\leq\;C_{3}\Bigg\{\sqrt{\kappa^{2}\sum_{k=1}^{3}\frac{p_{k}}{\sigma^{2}_{r_{k}}\!\big(\mathcal{M}_{k}(X^{*})\big)}}\;+\;\frac{\xi_{(r_{1},r_{2},r_{3})}}{\|X^{*}\|_{\mathrm{F}}}\Bigg\}.$ (8)
We next show that the rate in (7) is minimax optimal by proving a matching lower bound.
Theorem 4 (Minimax lower bound).
Let $r_{1},r_{2},r_{3},p_{1},p_{2},p_{3}$ be arbitrary integers such that $r_{i}\leq\min\{p_{i},\prod_{j\not=i}p_{j}\}$ for all $i\in\{1,2,3\}$. Let $\mathcal{T}_{(r_{1},r_{2},r_{3})}$ be the class of Tucker low-rank tensor as in (5) and for $\xi>0$, let
$\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}=\big\{B\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}:\inf_{A\in\mathcal{T}_{(r_{1},r_{2},r_{3})}}\|B-A\|_{\mathrm{F}}\leq\xi\big\}.$
Consider the model $Y=X^{\ast}+Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ where the entries of $Z$ are i.i.d. $\mathrm{subGaussian}(0,\kappa^{2})$. For any ranks $(r_{1},r_{2},r_{3})$ and any $0\leq\xi\leq\kappa^{2}p_{1}p_{2}p_{3}$, there exist a universal constant $c>0$ such that
$\inf_{\widehat{X}}\ \sup_{X^{\ast}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{\ast}\bigr\|_{\mathrm{F}}^{2}\$$\geq\ c\Bigl\{\kappa^{2}\big(\textstyle\sum_{k=1}^{3}p_{k}r_{k}+r_{1}r_{2}r_{3}\big)+\xi^{2}\Bigr\}.$ (9)
Tightness. The upper bound in Theorem 3 and the matching lower bound in Theorem 4 together show that the error bound in (7) is minimax optimal up to universal constants over the approximately Tucker low-rank class $\mathcal{T}^{\xi}_{(r_{1},r_{2},r_{3})}$.
Advantages of low-rank representations. Even when $X^{*}$ has full Tucker rank $(p_{1},p_{2},p_{3})$, it is common to approximate $X^{*}$ by a Tucker rank-$(r_{1},r_{2},r_{3})$ estimator. This is because (i) the approximation bias can be small when $X^{*}$ is well approximated by a low-rank tensor, so balancing bias and variance may reduce estimation error; and (ii) low-rank structure can yield substantial gains in computation and storage, as demonstrated in the following remark.
Remark 3 (Computational and storage savings from tensor low-rank estimation).
@@
$Y_{(r_{1},r_{2},r_{3})}=S\times_{1}U_{1}\times_{2}U_{2}\times_{3}U_{3}$
@@
$s=t_{2}^{\top}S^{\prime}\in\mathbb{R}^{r_{3}}\ (\text{cost }\approx 2r_{2}r_{3}),\quad w=s^{\top}t_{3}\in\mathbb{R}\ (\text{cost }\approx r_{3}).$